OpenStudy (anonymous):

HELPP! Let F be a function defined by f(x) 1-2sinx (x is less or equal to 0) and e^-4x (x is greater than 0) For x does not equal 0, express f '(x) as a piecewise defined function. Find the value of x for which f ' (x) = 3

OpenStudy (across):

\[f(x)=\begin{cases} 1-2\sin x & \text{ if } x\leq0 \\ e^{-4x} & \text{ if } x>0 \end{cases}\]Is this what you mean?

OpenStudy (across):

\[f'(x)=\begin{cases} -2\cos x & \text{ if } x\leq0 \\ -4e^{-4x} & \text{ if } x>0 \end{cases}\]

OpenStudy (anonymous):

yes the first one